The stopcock connecting a 1.00 L bulb containing argon gas at a pressure of 606 mm Hg, and a 1.00 L bulb containing carbon dioxide gas at a pressure of 672 mm Hg, is opened and the gases are allowed to mix. Assuming that the temperature remains constant, what is the final pressure in the system? mm Hg
Answer #1


We will use Boyles law
P1*V1 + P2*V2 = Pf*Vf
P1*V1 + P2*V2 = Pf*(V1+V2)
606*1.00 + 672*1.00 = Pf*(1.00+1.00)
1278 = Pf*2.00
Pf = 639 mmHg

Answer: 639 mmHg